OpenStudy (anonymous):

How would you simplify tan(sin^-1 (x/2)) to find its exact value? Is sin^2 theta+cos theta=2 true or false?

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle whose sine is \(\frac{x}{2}\) find the third side via pythagoras then you can find any trig ratio you want

OpenStudy (anonymous):

Does it matter that it's sin^-1?

OpenStudy (anonymous):

not sure what you are asking but i guess the answer is "of course it matters" if it was cosine it would be different

OpenStudy (anonymous):

did you find the third side?

OpenStudy (anonymous):

i think maybe i lost you there ?

OpenStudy (anonymous):

Sorry, I spent some time doing the calculation with identities instead of a triangle and ended up with \[ \tan(\sin^-1 x/2)= (x/2)/(√1-(x^2/4))=x/(2√(1-(x^2/4)))\] Would this result in the same answer as using the triangle?

OpenStudy (amilapsn):

yep